OpenStudy (anonymous):

The eighth term in a geometric sequence is 729. The first term in the sequence is 1/3. What is the fifth term? Please help and show work! I have no idea how to do this

OpenStudy (anonymous):

Do we multiply by 3? 1/3 3 0 27 81 243 729 ^ | | fifth term

OpenStudy (anonymous):

fist term = 1/3

OpenStudy (anonymous):

eight term is 729

OpenStudy (anonymous):

we know formula for GP tn = ar^(n-1)

OpenStudy (anonymous):

t1 = 1/3

OpenStudy (anonymous):

t8 = 729

OpenStudy (anonymous):

t8/t1 = ar^7/a = 729/(1/3)

OpenStudy (anonymous):

cancel out 'a'

OpenStudy (anonymous):

r^7 = 729/(1/3)

OpenStudy (anonymous):

r^7 = 729*3

OpenStudy (anonymous):

r^7 = 2187

OpenStudy (anonymous):

Answer is t5 = 27

OpenStudy (anonymous):

Can you explain how you set up the formula please?